OpenStudy (nottim):

If the velocities in such a beam followed the Maxwell-Boltzmann distribution the average velocity would be given by: < v >= a π ve−av2 v=0 ∞ ∫ dv where a = m / 2kT . What would be the average velocity in m/s of a beam of O2 molecules at 100 K?

OpenStudy (nottim):

\[<v>=\sqrt{\frac{ m/2KT }{ \pi }}\int\limits_{v=o}^{\infty}ve ^{-av ^{2}}dv\]

OpenStudy (nottim):

Getting kicked out my my working space soon. Will be back in an hour.

OpenStudy (nottim):

Hiyo

OpenStudy (empty):

So basically you want to find <v> by evaluating the stuff on the right. So there are two major things to do: 1) plug in all the values for m, k, and T. 2) solve the integral/look it up in a chart. I suggest plugging in everything once you're sure you have the integral solved.

OpenStudy (nottim):

This is a little embarassing, but I can't, for the life of me, remember how calculus works.

OpenStudy (aaronq):

If you don't want to revisit calculus (which i think you should), you can use wolfram alpha or other software. Look here, https://www.wolframalpha.com/input/?i=0+to+inf+integral+xe%5E(ax%5E2)+dx

OpenStudy (aaronq):

I omitted the stuff under the squared root sign since it's only a constant, btw.